5. A-4.0nc charge and a +5.Onc charge are separated by 50cm along the x-axis. Show that a -2 nC charge placed @ x = -4.24 m will experience zero net force. - 4.0 nC 5.0 nC 0.5m
Answer #1

Force on -2nC by - 4nC = kq1q2 / r1^2

= 9*10^9*4*10^-9*2*10^-9 / 4.24^2

= 4*10^-9 (to the left )

Force on -2 nC by 5 nC = 9*10^9*2*10^-9*5*10^-9 / (4.74)^2

= 4*10^-9 ( to the right )

so net force = 0
Answer #2
Three point charges are arranged on a line. Charge q3 = +5.00 nC and is at the origin. Charge q2 = -3.00 nC and is at x = +4.00 cm. Charge q1 is at x = +2.00 cm. What is q1 (magnitude and sign) if the net force on q3 is zero?